OpenStudy (anonymous):

Mario Agnello is opening a new pizza shop in town. He plans to offer a 12 -inch diameter pizza and an 18 -inch diameter pizza. He has set the price of a 12 -inch cheese pizza at 8 dollars , based on the amount of ingredients needed to cover that size crust. If Mario would like to keep the pricing proportional, what should he charge for an 18 -inch cheese pizza? Show and explain your work.

OpenStudy (anonymous):

$8 for 12 inch pizza half of it $4 for 6 inch pizza three times $12 for 18 inch pizza

OpenStudy (anonymous):

\[12/18=8/x\]\[12x=18(8)\]\[12x=144\]\[x=12\] so answer is 12 dollars